a) How much water will it take to fill the swimming pool with the dimensions 2.5m by 10m by 6m? ( 1m^3=1000L)
b) How long will it take to fill the pool at a rate of 5L every 30 seconds?

Found 2 solutions by checkley77, jojo14344:
Answer by checkley77(12844)

2.5*10*6=150 m^3 IS THE VOLUME OF THE POOL.
150*1000=150,000 L
150,000/5=30,000*30=900,000 SECONDS OR 8.33 HOURS TO FILL THE POOL.

Answer by jojo14344(1513)

There you go, since you've given the conversion, just get the volume, 2.5m*6m*10m=150cu.m.*(1000l/m^3) = 150,000liters
For the time to fill it up, divide by 5liters/30 sec=1liter/6sec,
150,000 liters
--------------
1 liter / 6sec
=900,000 sec x 1 min/60sec= 15,000 mins x 1hour/60min
= 250 hours x 1day/24 hours =10.41666667 days
= 10 days & 10 hours, time to fill the pool
Thank you,
Jojo